OpenStudy (yessenia27):

Solve x3 = 64 over 27.

OpenStudy (jakesaurus1127):

Is that supposed to say \[x^{3}\]

OpenStudy (haleyelizabeth2017):

\[x^3=\frac{64}{27}\]?

OpenStudy (yessenia27):

Yes

OpenStudy (xxstudyingstuffxx):

4 over 3

OpenStudy (haleyelizabeth2017):

Please do not give direct answers.

OpenStudy (sunnnystrong):

Just take the cube root of 64/27 The answer is 4/3

OpenStudy (xxstudyingstuffxx):

Sorry

OpenStudy (haleyelizabeth2017):

Don't be sorry (:

OpenStudy (xxstudyingstuffxx):

\[\frac{ 4 }{ 3 }\]

OpenStudy (yessenia27):

But I want direct answers😂

OpenStudy (yessenia27):

Solve x2 = −169.

OpenStudy (xxstudyingstuffxx):

☺☺☺☻☻☻☺☺☺☻☻☻☺☺☺ ☺☺☺☻☻☻☺☺☺☻☻☻☺☺☺ ☺☺☺☺☺☺☺☺☺☺☺☺☺☺☺ ☺☺☺☻☻☻☺☺☺☻☻☻☺☺☺ ☺☺☺☻☻☻☺☺☺☻☻☻☺☺☺ ☺☺☺☺☺☺☻☻☻☺☺☺☺☺☺ ☺☺☺☺☺☺☻☻☻☺☺☺☺☺☺

OpenStudy (yessenia27):

Solve x2 = −169.

OpenStudy (jakesaurus1127):

\[x^{3}=\frac{ 64 }{ 27 } \rightarrow \sqrt[3]{x^{3}}=\sqrt[3]{\frac{ 64 }{ 27 }} \rightarrow x=\frac{ \sqrt[3]{64} }{ \sqrt[3]{27} } \rightarrow x=\frac{ 4 }{ 3 }\]

OpenStudy (xxstudyingstuffxx):

▓▓▓ ▓▓▓ ▓▓▓ ▓▓▓ ▓▓▓ ▓▓▓ ▓▓▓ ▓▓▓ ▓▓▓ ▓▓▓ ▓▓▓ ▓▓▓ ▓▓▓▓▓▓ ▓▓▓▓▓▓

OpenStudy (plainntall):

You can use a calculator for direct answers. That isn't fun though. \[x ^{2}=-169\] is tricky because you can't multiply anything by itself and get a negative number. Mathematicians like to play pretend though so they made up and imaginary number and called it i \[i==-1\] That makes your problem \[x=(\sqrt{169})i\]

OpenStudy (yessenia27):

What's the answer